OpenStudy (anonymous):

Consider the following pseudocode function. function W(n is in Z) if n > 0 then if n is odd then return 2n else return n + 1 else return n − 1 Compute the values returned by the following function calls. (a) W(14) = (b) W(−20) = (c) W(9) =

OpenStudy (anonymous):

We see that if \(n\) is positive and odd we get back \(2n\); if it's even, we get back \(n+1\). If it's negative, we get back \(n-1\). Break it down now: (a) \(n=14\) is positive and even so we get back \(W(14)=14+1=15\) (b) \(n=-20\) is negative so we get back \(W(-20)=-20-1=-21\) (c) \(n=9\) is positive and odd so we get back \(W(9)=2(9)=18\)

OpenStudy (anonymous):

thank u @oldrin.bataku